If the measure of an angle is twice the measure of its complement, what is the measure of an angle?
Answer by mananth(16946)

let the angle be x
complement = 90-x
..
x=2(90-x)
x=180-2x
3x=180
x= 60
..
the angles are 60,30
....
m.ananth@hotmail.ca